# SuperGrad: a differentiable simulator for superconducting processors

### v2

@@ 5.1 Library design
Neural network libraries such as Flax [44] and Haiku [27] ensure flexible usage of the JAX framework, mitigating the conflict between OOP and JAX’s requirement for pure functions. In SuperGrad, we use Haiku to transform the quantum processor composited by objects into pure functions. Haiku can help us collect all parameters declared in a computation into a Python dictionary and pass the dictionary through a transformed pure function. In some quantum processor optimization tasks, it is reasonable to partition the parameters into trainable and non-trainable sets, and Haiku provides utilities for manipulating the parameters dictionary.
@@ 5.2 SuperGrad Enhancement Proposals
The usability of our current version SuperGrad 0.2.1 can be enhanced in several respects, particularly concerning the user’s interaction and the system’s flexibility:
@@
    We are considering to make SCGraph a JAX-differentiable data structure. The transformed pure functions will use SCGraph as input data, and the gradient of the objective function will take the same data structure as SCGraph. Moreover, the enhanced SCGraph will support functions for handling random deviations and parameter sharing, which further expands its functionality.
@@ 5.3 Realization of other superconducting qubit architectures
5.4 Migrating to the Tensor Network
@@ 7 Acknowledgement
The authors would like to thank Wangwei Lan for constructive discussions. ZAW and XW acknowledge the partial support from the National Key Research and Development Program of China (2021YFA1401902). XTN acknowledges the partial support by Guangdong Provincial Quantum Science Strategic Initiative (GDZX2200001).
@@ Appendix C Simultaneous Gate Simulation Details
Considering the presence of higher energy levels and adjacent qubits, the optimal frequency $\omega_{d}$ is derived by analyzing the composited quantum system. Furthermore, arbitrary $Z$-rotations are incorporated both before and after the gate operations. We also allow arbitrary $Z$-rotations before and after the gates since they can be essentially free [51].
@@ Appendix D Details of time evolution computation
The commonly used basis for multi-qubit processors is the product basis (the tensor product of single-qubit eigenbases) or the multi-qubit eigenbasis. This latter basis is derived by diagonalizing the idling Hamiltonian matrix in the product basis and then adjusting the phases of its eigenstates [17]. The product basis offers the advantage of storing the local Hamiltonian in the qubit subspace, facilitating efficient time evolution through the use of a series of 2-body operators. However, the local Hamiltonian in the multi-qubit eigenbasis tends to occupy the full Hilbert space, which significantly degrades the performance during time evolution.
Computing in different basis is almost equivalent, but some information will be lost when we truncate the matrices to the computational subspace. The multi-qubit eigenbasis is considered to be a good approach for the experimental basis [54]. We suggest a method that utilizes the multi-qubit eigenbasis to represent the final state while employing the product basis for time evolution calculations. The primary computational cost here involves the diagonalization of the idling Hamiltonian. The time complexity for diagonalization is $O(d^{3N})$, where $N$ represents the number of qudits and $d$ denotes the dimension of the qudit subsystem. This complexity becomes prohibitive for large quantum processors. Therefore, finding an efficient method to approximate the eigenstates of the idling Hamiltonian remains a question for future study. Alternatively, employing the product basis to represent the final state may reduce the diagonalization overhead but increase the error.
We begin the process by diagonalizing the idling Hamiltonian and use the eigenbasis for the time evolution unitary matrices. We denote the eigenstates by the matrix $V$, which is written in the product basis $B_{\textrm{product}}$. In this work, the Hamiltonians are always written in the product basis since we want to use trotter decomposition. Therefore, the time evolution unitary in the eigenbasis is
@@
That’s valuable to note that the dynamics of a density matrix could be characterized using a superoperator, which is computed by solving the vectorized Lindblad master equation [56]. In this way, SuperGrad migrated the Trotterization solver and backpropagation method to the open quantum systems.

### v3

@@ 5.1 Library design
Neural network libraries such as Flax [48] and Haiku [27] ensure flexible usage of the JAX framework, mitigating the conflict between OOP and JAX’s requirement for pure functions. In SuperGrad, we use Haiku to transform the quantum processor composed of objects into pure functions. Haiku can help us collect all parameters declared in a computation into a Python dictionary and pass the dictionary through a transformed pure function. In some quantum processor optimization tasks, it is reasonable to partition the parameters into trainable and non-trainable sets, and Haiku provides utilities for manipulating the parameters dictionary.
@@ 5.2 SuperGrad Enhancement Proposals
The usability of our current version SuperGrad 0.2.3 can be enhanced in several respects, particularly concerning the user’s interaction and the system’s flexibility:
@@
    We are considering making SCGraph a JAX-differentiable data structure. The transformed pure functions will use SCGraph as input data, and the gradient of the objective function will take the same data structure as SCGraph. Moreover, the enhanced SCGraph will support functions for handling random deviations and parameter sharing, which further expands its functionality.
@@
5.3 Role of SuperGrad in experiments
SuperGrad, like other simulation software, can be an invaluable tool in design optimization, cost-efficient development acceleration, pre-experimental planning, and post-experimental validation for fabricating superconducting processors, as is the case in other fields, such as aircraft, robotics, and semiconductor chip designs. While the exact simulation of a complex processor is generally unfeasible, the software complements expensive design iteration and device testing, and helps plan and understand the experimental measurements of device behavior. To showcase the multifaceted capabilities of SuperGrad, we provide the following two scenarios.
First, we often need to prototype novel processor designs with new types of qubits and couplings. For this purpose, we need to estimate important quantities such as qubit frequencies and 2-qubit gate fidelities. We then need to optimize the device design for fabrication. The experimental data obtained by measuring the fabricated devices can be employed to refine the simulation, progressively enhancing the predictive power of simulation models. As a result, the trial-and-error iterations can be substantially accelerated through this synergy between simulations and experiments.
Secondly, most of the time, we are looking for robust features in simulations and experiments. For example, in the experimental spectra data in Fig. 6, the transition frequencies $f_{01}$ can have small shifts due to two-level systems and other sources in the processors. However, for a working qubit, the overall shape of the spectra is close to the theoretical curve $f_{01}(\varphi_{\text{ext}})$, and therefore we can extract a good approximation of qubit parameters. Another example is the inaccuracy from control pulse distortions, which means that the qubits never experience the exact same control pulses as the simulation. However, we can use the simulation software to verify that for a gate scheme, the pulse parameters can be calibrated to compensate for pulse distortions. The average fidelities after calibration are thus a more robust quantity that can be extracted from simulation.
@@ 5.4 Realization of other superconducting qubit architectures
5.5 Adding a Tensor Network Backend
@@ 7 Acknowledgement
The authors would like to thank Wangwei Lan for constructive discussions. ZAW and XW acknowledge the partial support from the National Key Research and Development Program of China (2021YFA1401902). XTN acknowledges the partial support by Guangdong Provincial Quantum Science Strategic Initiative (GDZX2203001, GDZX2403001) and NSFC (No.92476206). FW and HHZ are supported by Zhongguancun Laboratory.
@@ Appendix C Simultaneous Gate Simulation Details
Considering the presence of higher energy levels and adjacent qubits, the optimal frequency $\omega_{d}$ is derived by analyzing the composited quantum system. Furthermore, arbitrary $Z$-rotations are incorporated both before and after the gate operations. We also allow arbitrary $Z$-rotations before and after the gates, since they can be essentially free [55].
@@ Appendix D Details of time evolution computation
The commonly used basis for multi-qubit processors is the product basis (the tensor product of single-qubit eigenbasis) or the multi-qubit eigenbasis. This latter basis is derived by diagonalizing the idling Hamiltonian matrix in the product basis and then adjusting the phases of its eigenstates [17]. The product basis offers the advantage of storing the local Hamiltonian in the qubit subspace, facilitating efficient time evolution through the use of a series of 2-body operators. However, the local Hamiltonian in the multi-qubit eigenbasis tends to occupy the full Hilbert space, which significantly degrades the performance during time evolution.
Computing in a different basis is almost equivalent, but some information will be lost when we truncate the matrices to the computational subspace. The multi-qubit eigenbasis is considered to be a good approach for the experimental basis [58]. We suggest a method that utilizes the multi-qubit eigenbasis to represent the final state while employing the product basis for time evolution calculations. The primary computational cost here involves the diagonalization of the idling Hamiltonian. The time complexity for diagonalization is $O(d^{3N})$, where $N$ represents the number of qudits and $d$ denotes the dimension of the qudit subsystem. This complexity becomes prohibitive for large quantum processors. Therefore, finding an efficient method to approximate the eigenstates of the idling Hamiltonian remains a question for future study. Alternatively, employing the product basis to represent the final state may reduce the diagonalization overhead but increase the error.
We begin the process by diagonalizing the idling Hamiltonian and using the eigenbasis for the time evolution unitary matrices. We denote the eigenstates by the matrix $V$, which is written in the product basis $B_{\textrm{product}}$. In this work, the Hamiltonians are always written in the product basis since we want to use Trotter decomposition. Therefore, the time evolution unitary in the eigenbasis is
@@
It’s valuable to note that the dynamics of a density matrix could be characterized using a superoperator, which is computed by solving the vectorized Lindblad master equation [60]. In this way, SuperGrad migrated the Trotterization solver and backpropagation method to the open quantum systems.